40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10297     Entry time: Wed Jul 30 11:15:44 2014
Author: Akhil 
Type: Update 
Category: General 
Subject: Weekly Update 

 Plan for the week:

  • PID loop design and testing with the Green laser beat note by actuating the arm cavity length.
  • Beat note readout on MEDM screens and Strip tool.
  • Calibration of the laser frequency response to PZT signal in MHz/V using a test DC input(Koji assigned me this task because this calibration has not been done and is very useful).

Inside the Lab:

  • Placing the FOL box sometime in the afternoon today(with supervision of Manasa / EricQ).
  • Calibration of the PZT(Today or tomorrow).

 

ELOG V3.1.3-